OpenStudy (anonymous):

Two trains A and B start from Delhi and Kolkata towards Kolkata and Delhi respectively. After passing each other, they take 10 hours and 2 hours 30 minutes respectively to reach their destinations. If Train A is moving at the speed of 37.5 kmh", what is the speed of Train B?

OpenStudy (kropot72):

Let s = speed of train B The distance travelled by train A after passing train B = 37.5 * 10 = 375 km Let t = the time taken before the trains pass. Distance travelled by train B before the trains pass = st =375 km Therefore s = 375/t .........................................(1) The closing speed of the trains = 37.5 + s Total distance travelled by trains before passing = 375 + 2.5s \[time=\frac{distance}{speed}\] \[t=\frac{375+2.5s}{37.5+s}\] which gives \[37.5t+st=375+2.5s .................(2)\] Substituting the value for s from (1) into (2) gives \[37.5t+375=375+\frac{2.5\times 375}{t}\] which simplifies to \[37.5t=\frac{2.5\times 375}{t}\ .........................(3)\] Rearranging (3) we get\[t=\sqrt{\frac{2.5\times 375}{37.5}}\] Speed of train B = 375/t